A camper is about to drink his morning coffee. He pours 400 grams of coffee, initially at 75.0°C, into a 250-g aluminum cup, initially at 16.0°C. What is the equilibrium temperature of the coffee-cup system, assuming no heat is lost to the surroundings? The specific heat of aluminum is 900 J/(kg·K). Assume that the specific heat of coffee is the same as the specific heat of water.
Answer #1

Let final temprature beT.

Heat lost by coffee= mCdT

                                =0.400*4180*(75.0-T)

Heat gained by cup= 0.250*900*(T-16)

For no heat lost.

Heat lost by coffee= Heat gained by cup

0.400*4180(75-T)=0.250*900(T-16)

T=68 ?